f is the real valued function defined by f(x) = 1 / (1 + e^x).
(i) Draw the variation table of f and sketch its representative curve.
(ii) Show that the point I(0, 1/2) is the center of symmetry of the curve of f.
1
Expert's answer
2021-01-27T10:21:30-0500

(i)




(ii)

By the definition, let "f(x)"  be a real-valued function of a real variable. Then "f(x)"  is odd if the following equation holds for all "x"  such that "x"  and "(-x)"  are in the domain of "f(x)" :



"f(x)=-f(-x)"

Geometrically, the graph of an odd function has rotational symmetry with respect to the origin.

(More information: https://en.wikipedia.org/wiki/Even_and_odd_functions)


To prove that the point "p.I(a,b)" is the center of symmetry of the function "f(x)" , it suffices to prove that the function "g(x)=f(x+a)-b" is odd.


In our case, we want to prove that point "p.I\left(0,1/2\right)" is the center of symmetry, so we need to consider the function



"f(x)=\frac{1}{1+e^x}\to g(x)=f(x+0)-\frac{1}{2}=\frac{1}{1+e^x}^{/2}-\frac{1}{2}^{/ \left(1+e^x\right)}=\\[0.3cm]
=\frac{2-1-e^x}{1+e^x}=\frac{1-e^x}{1+e^x}\to\boxed{g(x)=\frac{1-e^x}{1+e^x}}"

It remains to prove that the new function "g(x)" is odd.



"g(-x)=\frac{1-e^{-x}}{1+e^{-x}}=\frac{e^{-x}\left(e^x-1\right)}{e^{-x}\left(e^x+1\right)}=
\frac{-\left(1-e^x\right)}{1+e^x}=-g(x)\\[0.3cm]
\boxed{g(-x)=-g(x)\to g(x)-\text{odd function}}"

Conclusion,

Since "g(x)" turned out to be an odd function, it means that its center of symmetry is the origin of coordinates "O(0,0)" . This means that the function "f(x)" has a center of symmetry, point "I(0,1\/2)" .

Q.E.D.